OpenStudy (anonymous):

Solve for w, x, y, & z. w+x+y+z=18 w=x+y+z w+x-y+z=6 2w=3y 2x+y=8

OpenStudy (callisto):

w+x+y+z=18 -(1) w=x+y+z -(2) w+x-y+z=6 -(3) 2w=3y -(4) 2x+y=8 -(5) Sub. (2) into (1) to solve w. w+w=18 w = ...?

OpenStudy (callisto):

@wendy76 ?

OpenStudy (anonymous):

ok hold on

OpenStudy (kainui):

Better off learning linear algebra lol.

OpenStudy (callisto):

@wendy67 What have you got for w?

OpenStudy (anonymous):

ok wait sub 2 into 1? on what equation?

OpenStudy (anonymous):

@Kainui i'd be glad too but sadly i am taking math 2 but starting with math3 so we took a huge leap foward in math

OpenStudy (callisto):

w+x+y+z=18 -(1) w=x+y+z -(2) w+x-y+z=6 -(3) 2w=3y -(4) 2x+y=8 -(5) Actually, I labelled the equations with (number) So, Sub. (2) into (1) means sub the second equation into the first one. Since w=x+y+z (from the second equation) and w+x+y+z=18 (from the first equation) Sub x+y+z = w into the first equation and you'll get w + w = 18 Understand?

OpenStudy (anonymous):

oh im sorry didnt really understand what you meant by the #s on the side and yea i got it now so let me try then..

OpenStudy (callisto):

It's okay. I'm sorry... I should have explained it.

OpenStudy (anonymous):

No your fine, and i got 2x+2y+2z=18

OpenStudy (callisto):

Hmm..... How did you get that?

OpenStudy (anonymous):

ok so since w=x+y+z (2nd eq.) i plugged into the first: (x+y+z)+x+y+z=18 and got that^

OpenStudy (callisto):

Well.... You can do so... but that doesn't make it easier... Instead you can do it in this way... w=x+y+z (2nd eq.) That means x+y+z =w. Put it into the first one: w + (w) = 18.. Does it make sense to you?

OpenStudy (anonymous):

hmm does it mean that x,y,x got canceled and your left with w?

OpenStudy (anonymous):

*z

OpenStudy (callisto):

Not really, but substitute x+y+z = w into the equation..

OpenStudy (anonymous):

so if i substitude would it give me x+y+z+x+y+z=18?

OpenStudy (callisto):

If you sub w = x+y+z, yes. But you can do it in another way! Since x+y+z = w!

OpenStudy (anonymous):

ooooh i get it now so it WOULD be w+(w)=18 2w=18 w=9 ?

OpenStudy (callisto):

Yes!!!

OpenStudy (callisto):

Now, consider the 4th equation. 2w=3y Sub. w = 9 (which you've found) into the fourth equation. You can get the value of y. Can you do it?

OpenStudy (anonymous):

yes i already did all of them a got w=9 x=1 y=6 z=2

OpenStudy (anonymous):

*and

OpenStudy (callisto):

Yup :)

OpenStudy (anonymous):

ok well thank you (:

OpenStudy (callisto):

Welcome :)